### v2

# Broadband biphoton source for quantum optical coherence tomography based on a Michelson interferometer

Abstract
Broadband correlated photon pairs (biphotons) are valuable in quantum metrology, but current generation methods either involve complex nonlinear structures or lack sufficient bandwidth and brightness. In this work, we theoretically describe and experimentally demonstrate a novel technique for generation of a bright collinear biphoton field with a broad spectrum, achieved by using a tightly focused pump in a bulk nonlinear crystal. As the most straightforward application of the source, we employ Michelson interferometer-based quantum optical coherence tomography (QOCT). Utilizing the source enables the demonstration of record resolution and dispersion cancellation for this QOCT scheme.
preprint: APS/123-QED
I Introduction
Correlated photon pairs (biphotons) obtained through spontaneous parametric down-conversion (SPDC) [1] are one of the primary tools in quantum optics. The key role is played by time correlations between the photons of a pair, which become stronger the broader the spectrum of the biphoton field. A comprehensive review [2] details various implementations of broadband biphotons. Some are based on spatially homogeneous nonlinear media, where bandwidth is limited by phase matching and dispersion relations [3, 4, 5, 6, 7, 8, 9, 10]. Others use inhomogeneous media, where phase matching conditions vary, allowing broader bandwidth limited by spatial modulation range only [11, 12, 13, 14, 15, 16, 17, 18]. Achieving a biphoton field with a bandwidth up to 150-200 THz is feasible with these techniques. However, achieving higher values is challenging because, in all the mentioned approaches (with the exception of [8]), spectral broadening is inherently linked to a decrease in spectral intensity. In the best-case scenario, integral intensity may be conserved, but this is typically divergent from real experimental conditions.
Broadband biphotons find applications in diverse quantum metrological problems. Examples include quantum clock synchronization [19, 20, 21, 22, 23], two-photon microscopy [24, 25, 26], two-photon spectroscopy [27, 28, 29]. Another notable and straightforward application is quantum optical coherence tomography (QOCT) [30], which extends the capabilities of classical optical coherence tomography (OCT) [31]. Early OCT systems employed a broadband light source and a Michelson interferometer (MI), with one arm serving as a scanning reference and the other containing the test sample. Despite advancements in OCT’s accuracy through intricate experimental designs and sophisticated data analysis, challenges persist, particularly in addressing limited scanning depth and resolution for highly dispersive samples.
A quantum version of OCT utilizes entangled photon pairs and the Hong-Ou-Mandel effect. This method, being dispersion-tolerant, offers double the resolution, reaching record values [32]. The initial hurdle was the requirement for a bright broadband biphoton source, typically produced by non-collinear or type-II SPDC. However, further research showed that QOCT could also be achieved using a brighter, broader, and more easily adjustable collinear type-I SPDC source, coupled with an MI [33, 34, 35, 36].
In this work, we experimentally study for the first time a novel biphoton broadening technique based on tightly focusing the pump and target modes [37], and apply it to MI-based QOCT. Ideally, this broadening approach preserves spectral intensity, leading to an increase in integral intensity with a broader spectral bandwidth.
The article unfolds as follows. In Section II, we provide a theoretical description of the source’s operation and detail its experimental implementation. Section III delves into the theoretical and laboratory aspects of the interferometer designed for QOCT. Finally, in Section IV, we present and discuss the experimental results of QOCT experiments.
II Bright source of broadband biphoton field
II.1 Theory
Figure 1: Experimental setup. a) Broadband biphoton source: $L$ – crystal length, $W$ – modes waist, $\Theta_{\text{p}}$ – pump walk-off angle, L1, L2 – lenses. b) QOCT Michelson interferometer: BS – beam splitter, DM – longpass dichroic mirror splitting biphotons at 1000 nm, M1 – reference mirror, M2 – sample mirror, standing on the piezo-translator which is attached to a mechanical translator, D1, D2 – single-photon detectors (SPD), based on Si avalanche photodiodes (APD) Laser components COUNT NIR, D3 – SPD, based on InGaAs APD (MPD PDM–IR).
A comprehensive theoretical description of our biphoton source can be found in [37]. Similar to the listed techniques [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], we employ SPDC process as a biphoton source. SPDC involves the decay of pump (p) photons into pairs of signal (s) and idler (i) photons, satisfying energy conservation and phase-matching conditions:
$\omega_{\text{s}}+\omega_{\text{i}}-\omega_{\text{p}}$ $=0,$ (1)
$\vec{k}_{\text{s}}+\vec{k}_{\text{i}}-\vec{k}_{\text{p}}$ $=\vec{\Delta}_{k},$ (2)
where $\omega_{\text{p,s,i}}$ are frequencies, $k_{\text{p,s,i}}$ are wave vectors, and $\vec{\Delta}_{k}$ is a phase mismatch, typically $\lesssim$ the inverse nonlinear media length $1/L$.
Since our biphoton source interfaces with an interferometer, our focus is on the photon pair rate in a single spatial target mode, filterable by a single-mode fiber. We assume a Gaussian profile for the pump, signal, and idler modes propagating along the $z$-axis (top inset in Fig. 1a):
$\begin{gathered}\vec{E}_{j}(\vec{r},t)=\frac{1}{2}\left[\vec{E_{j}}g_{j}(\vec{r})\operatorname{e}^{-i\omega_{j}t}+\text{ c.c. }\right],\\ \text{where }\quad g_{j}(\vec{r})=\operatorname{e}^{ik_{j}z}\operatorname{e}^{-\frac{x^{2}+(y-\Theta_{j}z)^{2}}{W^{2}_{j}}},\quad j=\text{p, s, i}.\end{gathered}$ (3)
Here, $\vec{E_{j}}$ is a field amplitude, $\Theta_{j}$ is a walk-off angle, and $W_{j}$ is a beam waist. Diffraction terms are omitted under the assumption that the Rayleigh length is significantly larger than other longitudinal scales of the system. According to [38], the photon pair rate in the selected signal and idler modes can be calculated using a mode overlap integral:
$R=\frac{8d^{2}P}{\pi^{6}W^{2}_{\text{p}}W^{2}_{\text{s}}W^{2}_{\text{i}}c^{3}}\frac{\omega_{\text{s}}\omega_{\text{i}}}{n_{\text{p}}n_{\text{s}}n_{\text{i}}}\int_{-\infty}^{+\infty}\text{d}\omega\\ \times\left|\int_{-\infty}^{+\infty}\text{d}x\text{d}y\int_{-L/2}^{L/2}\text{d}zg_{\text{p}}(\vec{r})g_{\text{s}}^{*}(\vec{r})g_{\text{i}}^{*}(\vec{r})\right|^{2},$ (4)
where $d$ is an effective second-order nonlinear susceptibility, $n_{\text{p}},\ n_{\text{s}},\ n_{\text{i}}$ are refractive indices for the pump, signal, and idler modes, $c$ is the speed of light, and $P=c\epsilon_{0}n_{\text{p}}\pi W_{\text{p}}^{2}\left(\left|E_{\text{p}}\right|/2\right)^{2}$ is the pump power.
Considering a collinear e-oo type-I phase matching, i.e., $\Theta_{\text{s}}=\Theta_{\text{i}}=0$, and assuming equal waists for simplicity ($W_{\text{p}}=W_{\text{s}}=W_{\text{i}}=W$), we focus on a low-waist case where an effective overlap length $L_{\text{eff}}=W/\Theta_{\text{p}}\ll L$. Thus, in (4), we can integrate over $z$ from $\infty$- ∞ to +$\infty$+ ∞. After evaluating the integral, we obtain:
$R=\underbrace{\frac{4d^{2}P\omega_{\text{s}}\omega_{\text{i}}}{3\pi^{3}c^{3}\epsilon_{0}n_{\text{p}}n_{\text{s}}n_{\text{i}}\Theta_{\text{p}}^{2}}}_{S_{0}}\underbrace{\int\operatorname{e}^{-3\left(\Delta_{k}\left(\omega_{\text{s}}\right)L_{\text{eff}}/2\right)^{2}}\text{d}\omega_{\text{s}}}_{B_{\omega}}.$ (5)
Here, the factor $S_{0}$ represents the spectral coincidence rate, which is independent of the beam waist $W$ and crystal length $L$. The factor $B_{\omega}$ defines the integral spectral bandwidth and depends on the effective interaction length $L_{\text{eff}}$ instead of $L$. As a result, both the total rate $R$ and the bandwidth $B_{\omega}$ are proportional to $1/\sqrt{W}$.
Note that the spectral shape defined by the function $\operatorname{e}^{-3\left(\Delta_{k}\left(\omega_{\text{s}}\right)L_{\text{eff}}/2\right)^{2}}$ is almost Gaussian and can be tailored by the shaping of the pump and target modes. This technique can be used to increase the SPDC heralding efficiency and single-photon spectral purity instead of employing complicated and non-adaptive non-linear domain engineering [39, 40, 41].
For the BBO crystal pumped at 405 nm, as studied in [37] and used in our experiment, the spectral coincidence efficiency is $S_{0}/P=125\ \frac{\text{cps}}{\text{THz}\times\text{mW}}$, and the total spectral bandwidth is given by $B_{\omega}=\frac{\kappa}{\sqrt{W}}$, where $\kappa=2\pi\times 298\ \text{THz}\times\sqrt{\mu\text{m}}$.
Therefore, the primary idea is to reduce the interaction length $L_{\text{eff}}$ and compensate for a small rate through high focusing of the pump and target modes. A similar approach was demonstrated in [9], where authors used extremely short nonlinear media, achieving a significant biphoton rate due to high focusing. Conversely, another biphoton broadening technique using pump focusing [4] is based on a quite different approach and cannot provide as extensive broadening as our technique.
II.2 Implementation
The schematic representation of the source is presented in Fig. 1a. To achieve dispersion cancellation, as detailed in [42, 43], we employed a narrowband pump laser (Toptica DLC DL pro HP 405) with a wavelength of 405 nm and a bandwidth of approximately 100 kHz. The laser radiation passed through a single-mode fiber and was then tightly focused onto a 1 mm-thick nonlinear crystal (BBO type-I). The resulting collinear degenerate biphoton field was collimated using a parabolic mirror and passed through a pump-blocking 450 nm longpass filter. Subsequently, another parabolic mirror was utilized to couple the field into a single-mode polarization-maintaining photonic crystal fiber (NKT Photonics LMA-PM-10), chosen for its ability to guide and preserve the broadband spectrum. The biphoton field was then directed into the Michelson interferometer, described in the next section.
The alignment of the pump and target beams was meticulously optimized to achieve the maximum coincidence count rate, measured directly after the output of the fiber, reaching a final value of 3 kcps. The pump power, measured in front of the crystal, was 8 mW. Post-experiment, the crystal was removed, and the pump waist in free space was measured as $W_{p}=5.7~{}\upmu\rm m$.
III Interferometer
III.1 Theory
A comprehensive theoretical description of biphoton field interference in a Michelson interferometer (MI) can be found in [34, 35, 43]. Here, we will provide the main basis and results. Consider an input biphoton field in a single spatial mode with the following spectral quantum state:
$\left|{{\Psi}_{\text{in}}}\right\rangle=\frac{1}{\sqrt{2}}\iint{\text{d}{{\omega}_{\text{s}}}\text{d}{{\omega}_{\text{i}}}}f\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)\hat{a}_{\text{in}}^{\dagger}\left({{\omega}_{\text{s}}}\right)\hat{a}_{\text{in}}^{\dagger}\left({{\omega}_{\text{i}}}\right)\left|\text{vac}\right\rangle.$ (6)
where $f(\omega_{\text{s}},\omega_{\text{i}})$ is a spectral amplitude, approximated as a double-Gaussian function with the two-photon spectral distribution [44, 45] (Fig. 2):
${{\left|f({{\omega}_{\text{s}}},{{\omega}_{\text{i}}})\right|}^{2}}=\frac{1}{\pi\delta\Delta}\underbrace{\operatorname{e}^{-\frac{{{\left({{\omega}_{1}}-{{\omega}_{2}}\right)}^{2}}}{2{{\Delta}^{2}}}}}_{\text{Phasematching}}\underbrace{\operatorname{e}^{-\frac{{{\left({{\omega}_{1}}+{{\omega}_{2}}-2{{\omega}_{0}}\right)}^{2}}}{2{{\delta}^{2}}}}}_{\text{Pump}}.$ (7)
Here, $\omega_{0}\equiv\omega_{\text{p}}/2$, $\delta$ is the standard deviation (StD) of the pump spectral distribution, and $\Delta=B_{\omega}/\sqrt{2\pi}$ is related to the phase matching bandwidth.
Now, let this field pass through an MI with a delay $T$ in a fixed arm and delay $\tau$ in a scanning arm (see Fig. 1b and Fig. S1 in Supplementary [43]). The MI output is split into two paths ending with single-photon detectors. Measuring the coincidence count rate depending on the MI delay $\tau$ allows obtaining an interferogram:
Figure 2: Double-Gaussian two-photon spectral distributions |$\absolutevalue{f(\omega_{\text{s}},\omega_{\text{i}})}$ |.
$M(\tau)=\frac{1}{16}\left[M_{\text{c}}+M_{0}(\tau)-M_{1}(\tau)+M_{2}(\tau)\right],$ (8)
$\text{where}\quad{{M}_{c}}$ $=\left(1+R\right)^{2},$ (9a)
${{M}_{0}}(\tau)$ $=2R\operatorname{e}^{-\frac{{{\left(T-\tau\right)}^{2}}}{2{{\Delta}^{-2}}}},$ (9b)
${{M}_{1}}(\tau)$ $=4r\left(1+R\right)\operatorname{e}^{-\frac{{{\left(T-\tau\right)}^{2}}}{8\Delta_{+}^{-2}}}\cos\left[{{\omega}_{0}}\left(T-\tau\right)\right],$ (9c)
${{M}_{2}}(\tau)$ $=2R\ \operatorname{e}^{-\frac{{{\left(T-\tau\right)}^{2}}}{2{{\delta}^{-2}}}}\cos\left[2{{\omega}_{0}}\left(T-\tau\right)\right].$ (9d)
Here $\Delta_{+}^{2}\equiv\delta^{2}+\Delta^{2}\approx\Delta^{2}$ and $r^{2}\equiv R$ is a sample reflectivity.
The total interferogram $M(\tau)$ is plotted in Fig. 3a. Its spectrum |$\absolutevalue{\tilde{M}(\omega)}$ | in Fig. 3b contains three peaks: $\tilde{M}_{0}$ centered at $\omega=0$ with StD $\Delta$, $\tilde{M}_{1}$ centered at $\omega_{0}$ with StD $\Delta/2$, and $\tilde{M}_{2}$ centered at $2\omega_{0}$ with StD $\delta$. The corresponding interferogram terms are plotted in Fig. 3c. Term $M_{0}$ represents a peak with StD $1/\Delta$, corresponding to HOM-interference. Term $M_{1}$ has twice the StD and corresponds to single-photon interference. Finally, $M_{2}$ term with a $1/\delta$ StD is equivalent to pump interference in an MI.
Figure 3: Total interferogram $M(\tau)$ (a), it’s spectrum (b) with terms $\tilde{M}_{0}(\omega)$, $\tilde{M}_{1}(\omega)$, $\tilde{M}_{2}(\omega)$ and corresponding interferogram terms $M_{0}(\tau)$, $M_{1}(\tau)$, $M_{2}(\tau)$ (c).
The filtering of the $M_{0}$ term allows achieving a resolution that is twice as good as that of the $M_{1}$ term, corresponding to standard OCT. Moreover, similar to the usual HOM-based QOCT, the $M_{0}$ term provides dispersion cancellation (see Supplementary [43] for details). However, we observe that terms $M_{0}$ and $M_{1}$ can be separated only if $\Delta\ll\omega_{\text{p}}/3$. Otherwise, the spectra $\tilde{M}_{0}(\omega)$ and $\tilde{M}_{1}(\omega)$ overlap, and the terms cannot be well separated anymore.
III.2 Implementation
For the basic experiment in QOCT, we constructed the setup illustrated in Fig. 1b. Biphoton pairs were emitted from the fiber, collimated using a parabolic mirror. The beam was then split on a non-polarized beam splitter (BS) into two channels: the sample channel and the reference channel. The sample channel included a mirror and an optional 2-mm wide glass plate to introduce dispersion. The reference channel consisted of a mirror mounted on a combined piezo-mechanical translator stage. The reflected beams were recombined on the subsequent BS.
To broaden the detection bandwidth and cover both the visible (VIS) and infrared (IR) ranges, we employed a longpass dichroic mirror (DM) with a cut-on wavelength $\lambda_{\text{c}}=1000~{}\text{nm}$. In both channels, biphotons were coupled with single-mode fibers using parabolic mirrors. The visible part was further split by a fiber beam splitter and directed to two Si-based avalanche photodiode (APD) detectors (Laser Components COUNT NIR), labeled as D1 and D2. The infrared part was directed to an InGaAs APD detector (MPD PDM-IR), labeled as D3. All detector outputs were connected to coincidence circuits, enabling coincidence measurements between the two visible-range detectors D1&D2 (VIS-VIS) and between the IR detector and both visible-range detectors: D3&D1 and D3&D2 (IR-VIS).
IV Experimental results and discussion
Initially, we conducted a $z$-scan using a sample consisting only of a mirror. The results are depicted in Fig. 4a – c. In Fig. 4a, we show raw single count rate interferograms for the VIS and IR channels (red and brown curves). A small shift between their envelopes indicates slight dispersion in either the sample or reference arm. Raw interferograms for VIS-VIS and IR-VIS coincidences are plotted in Fig. 4b, with the red and brown curves corresponding to each.
The absolute values of their Fast-Fourier Transform (FFT) spectra are shown in Fig. 4c as thin curves with the same colors. The raw spectra qualitatively match the theory presented in Fig. 3b (see the detailed analysis in Supplementary [43]). They exhibit three peaks corresponding to the terms $\tilde{M}_{0}$, $\tilde{M}_{1}$, and $\tilde{M}_{2}$, but due to the high bandwidth, the $\tilde{M}_{0}$ and $\tilde{M}_{1}$ peaks are partially overlapped. To extract the $\tilde{M}_{0}$ term, we used the spectral range $30-\omega_{p}/$ 3, and the range $\omega_{p}/3-0.75\omega_{p}$ for extracting the $\tilde{M}_{1}$ term. To account for the spectral dependencies of the detectors’ quantum efficiencies, $\eta_{\text{VIS}}(\omega)$ and $\eta_{\text{IR}}(\omega)$, we divided the raw spectra by the corresponding efficiency products (see details in Supplementary [43]).
The processed spectrum for the $\tilde{M}_{1}$ term, combined from both VIS-VIS and IR-VIS data, is plotted as a bold dotted curve. Its Gaussian fit, shown as a gray dashed curve, has a bandwidth $\text{FWHM}_{\omega}=2\pi\times 136~{}\text{THz}$, corresponding to a wavelength range of approximately 684–992 nm. This bandwidth is close to the estimated theoretical value of $\sqrt{\frac{4\log 2}{\pi}}B_{\omega}=2\pi\times 117$ THz (5), calculated for the measured pump waist $W=5.7\ \rm\mu m$.
Considering the spectral efficiency of the detectors and the bandwidth of coincidence detection (detailed information is provided in the Supplementary [43]), the registered coincidence count rate of 3 kcps suggests a total source generation rate of $2.7\ \frac{\text{kcps}}{\text{mW}}$. The source’s spectral generation rate of $S_{0}\approx 19~{}\frac{\text{cps}}{\text{THz}\times\text{mW}}$ is about 6.6 times smaller than the theoretical value of $125~{}\frac{\text{cps}}{\text{THz}\times\text{mW}}$ [37]. The mismatch between the experimental and theoretical results can be explained by various aberrations and the wavelength dependency of the target mode waist, which were not accounted for in the theory. However, this efficiency was sufficient to achieve a registered coincidence count rate of about 25–30 $\frac{\text{cps}}{\text{mW}}$ in the raw interferogram, whereas previous works reported results of 2–4 $\frac{\text{cps}}{\text{mW}}$ for similar QOCT experiments [35, 46].
Figure 4: Experimental results. a) Interferograms in single count rate; the red curve refers to the raw data for wavelengths $400-1000$ nm (VIS), registered by Si APD detectors, the brown curve refers to the raw data for wavelengths $1000-1500$ nm (IR), registered by InGaAs APD detectors, the blue curve refers to the processed data where detection efficiency is taken into account and signals are summed, and the grey dotted curve corresponds to the Gaussian approximation of the processed data. b) Interferograms in coincidence count rate; the red curve refers to the coincidences between two Si APDs (VIS-VIS), the brown curve shows coincidences between Si and InGaAs APDs (IR-VIS), the blue curve shows the processed data, and the grey dashed curve corresponds to the Gaussian fit. c) Coincidence spectra (absolute values of the Fast-Fourier Transform, FFT): the red thin line represents the raw VIS-VIS coincidence spectra, the brown thin line shows the spectrum of the raw IR-VIS coincidences, the bold red and brown solid lines correspond to the processed data used to extract the HOM term $M_{0}$, the dotted red and brown lines represent the processed spectrum used to extract the $\tilde{M}_{1}$ term, and the grey dashed curve shows its Gaussian fit. d) Interferograms in single count rate with dispersive media; the red curve refers to the raw VIS data, the brown curve refers to the raw IR data, and the blue curve refers to the processed data. e) Interferograms in coincidence count rate with dispersive media; the red curve refers to the raw VIS-VIS data, the brown curve refers to the raw IR-VIS data, the blue curve refers to the processed data, and the grey dashed line refers to the Gaussian fit.
Similarly, we divided the coincidence spectra related to the $\tilde{M}_{0}$ term by the products of quantum efficiencies. The result is shown in Fig. 4c as thick solid curves. Its inverse FFT, corresponding to the HOM peak with an additional sine term $M_{1}$, is shown in Fig. 4b as a blue curve. The FWHM of its Gaussian fit (gray dashed curve), which defines the QOCT resolution, was estimated to be approximately $0.46~{}\upmu\rm m$, slightly better than the record QOCT resolution of $0.54~{}\upmu\rm m$ reported in [46]. This value is consistent with the value $2\log(2)c/\text{FWHM}_{\omega}=0.49\ \upmu$m, corresponding to the measured spectral bandwidth.
We processed the single count rate interferograms in a similar manner, and the results are plotted in Fig. 4a as a blue curve. The FWHM of its Gaussian envelope (gray dashed curve), which defines the classical OCT resolution, was estimated to be $1.19~{}\upmu\rm m$. This is more than twice the QOCT resolution due to the presence of slight dispersion.
As a second step, we conducted a $z$-scan using a sample consisting of a mirror and a 2-mm glass plate, which introduces dispersion. The results are depicted in Fig. 4d–e. The single-photon count rate interferograms (Fig. 4d) show well-separated IR and VIS terms due to dispersion. The dispersion-induced broadening was significant enough that a single scan couldn’t cover the interference pattern range, requiring two scans to estimate the bandwidth ($\approx 50~{}\upmu\rm m$).
The raw coincidence interferograms and the filtered HOM term, accounting for detector efficiencies, are shown in Fig. 4e. The obtained resolution under such significant dispersion was $0.94~{}\upmu\rm m$, which is 50 times better than the OCT resolution but still twice as large as the QOCT resolution without dispersion. The non-ideal dispersion cancellation can be attributed to the influence of higher-order dispersion.
V Conclusion
We have introduced a cost-effective technique for generating a bright (2.7 kcps/mW pair generation rate) and broadband (136 THz bandwidth) biphoton field using a bulk nonlinear crystal with tight pump beam focusing. Demonstrating both theoretical insights and experimental validation, we applied this technique to quantum optical coherence tomography (QOCT), achieving an impressive axial resolution of $0.46\ \upmu$m, comparable to the record results reported in [32]. Significantly, we have demonstrated for the first time that Michelson interferometer-based QOCT exhibits dispersion cancellation, analogous to that observed in traditional Hong-Ou-Mandel-based implementations, resulting in at least a 50-fold improvement in resolution compared to classical Optical Coherence Tomography technique. Furthermore, our results suggest that the achieved QOCT resolution is limited by interferogram filtering procedures, indicating potential for further enhancement through the exploration of advanced filtering methods or modified interference schemes.